(1 point) Overweight baggage. Suppose weights of the checked baggage of airline passengers follow a nearly normal distribution with mean 48 pounds and standard deviation 5.2 pounds. Most airlines charge a fee for baggage that weigh in excess of 50 pounds. Determine what percent of airline passengers incur this fee. Round your answer to four decimal places. %
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 48
std deviation   =σ= 5.200

percent of airline passengers incur this fee:

probability =P(X>50)=P(Z>(50-48)/5.2)=P(Z>0.38)=1-P(Z<0.38)=1-0.648=0.3520~ 35.20%

(please try 35.03% if this comes wrong)